Not sure on how to solve this.
Use factoring to solve each equation for real values of the variable.
Please help me solve this problem: x - 13x^1/2 + 12 = 0

Answer by checkley77(12844)

x-13x^1/2+12=0
(SQRTX-12)(SQRTX-1)=0
SQRTX=12 SQUARE BOTH SIDES.
X=12^2
X=144 ANSWER.
SQRTX-1=0
SQRTX=1 SQUARE BOTH SIDES.
X=1^2
X=1 ANSWER.